OpenStudy (anonymous):

solve for x. 5/6+1/5x=7/10x-2/3

hero (hero):

1. Put like terms on the same side: \[\frac{5}{6} + \frac{2}{3} = \frac{7x}{10} - \frac{x}{5}\]

hero (hero):

Do you know how to combine fractions?

hero (hero):

You should know how to add \[\frac{5}{6} + \frac{2}{3}\]

OpenStudy (anonymous):

yes i do know how to do that

hero (hero):

Okay, well, what did you get after adding?

OpenStudy (anonymous):

3/2

hero (hero):

Okay good

hero (hero):

\[\frac{3}{2} = \frac{7x}{10} - \frac{x}{5}\]

hero (hero):

Now all you have to do is add the other side.

hero (hero):

You can combine the fractions if they have the same denominator, so multiply x/5 by 2/2

OpenStudy (anonymous):

2/10?

jimthompson5910 (jim_thompson5910):

What is the LCD in this case

OpenStudy (anonymous):

10?

hero (hero):

You can multiply by both sides by the LCD. That's on way to do it. I don't assume that the user knows how to find an LCD. But from now on, I think I'll just ask first.

hero (hero):

If they don't know, then I'll just proceed with my methods.

jimthompson5910 (jim_thompson5910):

I find it easier to get rid of the fractions despite whether they know the LCD or not

jimthompson5910 (jim_thompson5910):

I'll brb

hero (hero):

That's just the method you're used to @jim_thompson5910 . That's why you find it easier.

OpenStudy (anonymous):

ok Jim and Hero your methods confuse me btw

hero (hero):

Adding fractions is confusing? Why?

OpenStudy (anonymous):

cause i suck at math thats y and its hard to remember stuff from algebra 1 which i had 2 yrs ago thats y

hero (hero):

Well, you're lucky I'm not your teacher.

OpenStudy (anonymous):

well ya but we prob live on other sides of the U.S so you wouldnt have me

hero (hero):

Just because I use different methods doesn't mean I live on the other side of the country. Where are you from?

OpenStudy (anonymous):

Fl

hero (hero):

I lived in FL for 3 years. I'm familiar with the curriculum there.

OpenStudy (anonymous):

ah